### v3

@@ 5 Conclusions
This research comprehensively analyzes anthropometric measurements and their correlation with furniture dimensions in university settings. It proposes new dimensions for non-adjustable chairs and tables, as well as adjustable chairs and non-adjustable tables, to better suit the needs of students. The ANOVA test results demonstrated significant differences between the proposed and existing furniture dimensions, indicating the potential benefits of new measurements regarding comfort and ergonomics. The proposed dimensions showed reduced mismatch percentages for both males and females compared with existing furniture, suggesting a higher level of compatibility. This could improve comfort levels and lower the risk of MSDs among students who use the furniture. These findings highlight the importance of considering anthropometric measurements when designing furniture for educational environments.
The limitations of this research include the reliance on a specific population (university students), the exclusion of certain anthropometric measurements, and the lack of consideration of other factors influencing furniture design (e.g., cultural preferences and environmental constraints). Additionally, the proposed furniture dimensions may not be universally applicable and require further validation in other settings or populations. Future research could explore the implementation of the proposed dimensions in real-world classroom settings and assess their long-term impacts on student comfort and well-being. Additionally, further studies should investigate the effects of different furniture designs on student performance and productivity in academic settings.
@@
This study provides a comprehensive analysis of the ergonomic dimensions of furniture, particularly chairs and tables, in an educational setting. This highlights the importance of designing furniture that caters to anthropometric measurements of university students, ensuring comfort and promoting a healthy posture. By collecting anthropometric data and correlating them with existing furniture dimensions, this study identifies mismatches and provides recommendations for optimized furniture design. The findings have significant practical implications, as they can guide the development of furniture that promotes better ergonomics and reduces the risk of MSDs among students. Moreover, this study’s methodology and insights can be applied beyond the educational setting, informing furniture design for various industries and environments. This underscores the importance of considering cultural preferences and environmental constraints when designing furniture. This study contributes to the growing body of ergonomics research and can potentially improve the well-being and productivity of individuals in diverse settings.

### v4

@@ 6 Conclusions
This study underscores the critical importance of ergonomics in university computer laboratories, where students’ prolonged interactions with computers necessitate optimal ergonomic conditions. By proposing anthropometric-based furniture dimensions specifically tailored for university students, we aim to enhance the ergonomic design of these environments. Significant disparities were identified through meticulous analysis of 11 key anthropometric measurements and comparison with existing furniture dimensions. Statistical validation through ANOVA reinforced the credibility of our proposed dimensions, showing reduced mismatch percentages compared to existing furniture. This indicates higher compatibility and potential improvements in student comfort, as well as a reduction in the risk of MSDs. Our study also revealed significant disparities between existing furniture dimensions and students’ anthropometric data, particularly in crucial areas such as SW, BH, and STC. By identifying areas for improvement and proposing new dimensions aligned with students’ anthropometric measurements and international standards, our research offers practical insights that can inform future furniture design efforts and contribute to developing ergonomic standards in educational institutions.
Implementing these proposed dimensions can substantially improve the ergonomic design of university computer laboratories, ultimately promoting student health, well-being, and academic success. Moreover, this study has significant practical implications for educational institutions, furniture manufacturers, and students alike. Investing in furniture tailored to students’ body dimensions can create a more conducive learning environment, improving posture, reducing the risk of MSDs, and enhancing overall academic performance. By advocating for the adoption of ergonomically designed furniture, our research seeks to prioritize student welfare and productivity in university settings.
The contributions of the proposed work are as follows:
  1. 1.
  2. 2.
    By identifying the mismatch between existing furniture dimensions and anthropometric measurements, this study offers practical insights for designing ergonomic furniture in university environments and for improving comfort, health, and productivity among students and staff.
  3. 3.
  4. 4.
    We collected and analyzed anthropometric data from university students to provide valuable insights into the body measurements of the target population.
  5. 5.
    By comparing the mismatch between two different types of existing furniture and anthropometric data, this study offers a nuanced understanding of how different furniture designs affect users’ comfort and well-being.
@@
This study also provides recommendations for optimized furniture design. The findings have significant practical implications, as they can guide the development of furniture that promotes better ergonomics and reduces the risk of MSDs among students. Moreover, this study’s methodology and insights can be applied beyond the educational setting to inform furniture design for various industries and environments. This underscores the importance of considering cultural preferences and environmental constraints when designing furniture products. This study contributes to the growing body of ergonomics research and can potentially improve the well-being and productivity of individuals in diverse settings.
Ethical Approval
The Office of the Director for Research & Extension reviewed and approved this study under the approval code KUET/DRE/2023/15(6). Written informed consent was obtained from all participants prior to their inclusion in the study.
Data Availability Statement
The dataset used in this research, entitled “Anthropometric Data of KUET Students,” is available for further exploration and can be accessed on Mendeley [33]. It can be found at the following DOI link: https://dx.doi.org/10.17632/kw7fd465v7.2.
Funding
No financial support was provided by public, commercial, or non-profit funding bodies for this research.
CRediT authorship contribution statement
Anik Kumar Saha: Conceptualization, Data curation, Formal analysis, Investigation, Methodology, Software, Writing - original draft, Visualization. Md Abrar Jahin: Conceptualization, Data curation, Formal analysis, Investigation, Methodology, Software, Writing - original draft, Visualization. Md. Rafiquzzaman: Supervision, Validation, Writing - review & editing. M. F. Mridha: Validation, Writing - review & editing.
@@ Appendix A NMQ Survey Questionnaire Format
Appendix B Supplementary Tables and Figures
Table A1: Descriptive statistics of the anthropometric measurements taken for both male and female participants (in mm)
Measurements Gender Min Max Mean Standard Deviation Percentile
5th 50th 95th
PH Male 414 489 444 13 423 445 466
Female 373 447 415 14 395 414 438
SEH Male 188 308 234 17 205 235 261
Female 191 282 231 17 200 230 257
BPL Male 402 498 454 18 422 454 483
Female 394 478 447 17 412 448 471
AL Male 331 394 364 11 346 364 382
Female 322 366 342 10 327 342 356
TT Male 131 178 156 8 143 156 169
Female 126 175 145 10 129 145 161
SSH Male 470 556 512 15 489 512 539
Female 447 522 488 13 467 489 509
SCH Male 484 540 512 10 495 511 528
Female 469 524 494 11 475 494 510
EFL Male 426 471 450 8 437 450 462
Female 385 428 407 10 392 408 421
SEB Male 396 492 447 15 424 448 470
Female 390 464 423 17 396 421 453
HB Male 328 382 351 9 335 350 365
Female 344 390 366 8 353 367 379
BKL Male 496 548 521 10 505 522 536
Female 475 546 509 14 489 509 530
(a) Distribution of male anthropometric data
(b) Distribution of female anthropometric data
Table A2: ANOVA test results for non-adjustable furniture dimensions
5th 50th 95th
SH vs PH Male Observed 423 445 466 0.71 0.447 Accept
Expected 444 457 471
Female Observed 395 415 438 7.83 0.049 Reject
Expected 444 457 471
SW vs HB Male Observed 335 350 365 11.19 0.029 Reject
Expected 377 394 410
Female Observed 353 367 379 4.51 0.101 Accept
Expected 376 394 411
SD vs BPL Male Observed 423 454 484 5.38 0.081 Accept
Expected 389 406 424
Female Observed 413 448 471 3.61 0.13 Accept
Expected 389 406 424
BH vs SSH Male Observed 489 512 539 173.83 0 Reject
Expected 295 305 315
Female Observed 467 489 509 184.73 0 Reject
Expected 295 305 315
BW vs HB Male Observed 335 350 365 0.28 0.625 Accept
Expected 347 356 364
Female Observed 353 367 379 1.48 0.29 Accept
Expected 347 356 364
UEB vs SCH Male Observed 495 511 528 70.93 0.001 Reject
Expected 392 406 420
Female Observed 475 494 510 44.56 0.003 Reject
Expected 392 406 420
STH vs SEH Male Observed 206 235 261 0.17 0.694 Accept
Expected 231 241 252
Female Observed 201 230 258 0.45 0.538 Accept
Expected 231 241 252
STC vs TT Male Observed 143 156 169 58.13 0.002 Reject
Expected 80 89 97
Female Observed 129 145 161 28.58 0.006 Reject
Expected 80 89 97
TL vs BKL Male Observed 505 522 536 13.8 0.021 Reject
Expected 474 483 492
Female Observed 489 509 530 4.37 0.105 Accept
Expected 474 483 492
Table A3: ANOVA test results for adjustable chair with non-adjustable table dimensions
5th 50th 95th
SH (lowest limit) vs PH Male Observed 423 445 466 0.527 0.508 Accept
Expected 410 432 453
Female Observed 395 415 438 0.796 0.423 Accept
Expected 410 432 453
SH (highest limit) vs PH Male Observed 423 445 466 19.804 0.011 Reject
Expected 507 533 560
Female Observed 395 415 438 35.008 0.004 Reject
Expected 507 533 560
SW vs HB Male Observed 335 350 365 45.836 0.002  Reject
Expected 434 457 480
Female Observed 353 367 379 35.436 0.004 Reject
Expected 434 457 480
SD vs BPL Male Observed 423 454 484 1.023 0.369 Accept
Expected 410 432 453
Female Observed 413 448 471 0.337 0.593 Accept
Expected 410 432 453
BH vs SSH Male Observed 489 512 539 148.401 0 Reject
Expected 290 305 320
Female Observed 467 489 509 149.574 0 Reject
Expected 290 305 320
BW vs HB Male Observed 335 350 365 9.223 0.039 Reject
Expected 374 394 413
Female Observed 353 367 379 3.926 0.119 Accept
Expected 374 394 413
UEB vs SCH Male Observed 495 511 528 48.612 0.002 Reject
Expected 386 406 427
Female Observed 475 494 510 31.274 0.005 Reject
Expected 386 406 427
STH (lowest limit) vs SEH Male Observed 206 235 261 0.101 0.766 Accept
Expected 217 229 240
Female Observed 201 230 258 0.003 0.962 Accept
Expected 217 229 240
STH (highest limit) vs SEH Male Observed 206 235 261 26.745 0.007 Reject
Expected 314 330 347
Female Observed 201 230 258 28.057 0.006 Reject
Expected 314 330 347
STC (lowest limit) vs TT Male Observed 143 156 169 61.03 0.001 Reject
Expected 90 95 100
Female Observed 129 145 161 26.571 0.007 Reject
Expected 90 95 100
STC (highest limit) vs TT Male Observed 143 156 169 19.526 0.012  Reject
Expected 187 197 207
Female Observed 129 145 161 22.616 0.009 Reject
Expected 187 197 207
TL vs BKL Male Observed 505 522 536 16.011 0.016 Reject
Expected 434 457 480
Female Observed 489 509 530 8.739 0.042 Reject
Expected 434 457 480